# Accurate and thermodynamically consistent hydrogen equation of state for planetary modeling with flow matching

### v1

@@ Appendix E Effect of thermodynamic inconsistency on entropy calculation
The subtleties involved in the implementation of TI arise not only from different choices of integration path, but also the reference point. In fact, the latter can be seen as an alternative manifestation of the former, which can be understood by noting that an integration path can belong to different types (i.e., isotherm or isochore first) when viewed by different reference points. As an illustration, Fig. S5 also shows the entropy reported by Miguel et al. [4] (dashed lines). According to the authors, they processed the original REOS3 data by integrating along isotherm first, hence in accordance with the red solid lines of our result. However, these two sets of entropies clearly differ by an offset in the ab initio region of the main panel. The remedy of such an apparent disagreement turns out to be surprisingly easy: notice all the entropy isochores already align well, we just need to locate the reference point on another “correct” isotherm, say, 15000K. After doing this, the red solid lines then coincide perfectly with the dashed lines from Ref. [4], as shown in Figure S6.
Our reproduction of the entropy results by Miguel et al. [4] based on a specific integration path and reference point, as done above, demonstrates how fragile the “global” TI approach can be for an EOS with significant thermodynamic inconsistencies. Notice the solid and dashed lines in Fig. S5 differ by roughly , which can make a big difference on planetary model predictions as shown in the main text.
@@ Appendix F More details on the thermodynamically consistent construction of our EOS
To construct an EOS that is thermodynamically consistent across several distinct theories, such as ab initio DFT and SCvH, one needs to first calculate the free energy and entropy of each region separately using the standard TI approach. We then perform a two-dimensional spline interpolation over each set of free energy data, following a similar approach to Militzer and Hubbard [12]. This in principle allows us to reproduce the original pressure and energy data by performing partial derivatives. The quality of such reproduction is, of course, closely related to the original level of thermodynamic consistency in each region. Once the free energy interpolation for each separate region is ready, we can then glue them together by performing another spline interpolation along their boundaries. Note one can utilize relevant partial derivative values derived from the free energy interpolations on both sides to ensure a smoother connection.
@@
In principle, the free energy interpolation would be perfectly smooth if the absolute entropy and energy (measured relative to the same baseline) were exact on both sides. However, this is not the case in practice. First, it is generally believed that only the relative energies produced by DFT (using any given exchange-correlation functional) are reliable, rather than the absolute ones. This indicates that we are allowed to globally shift the DFT energies as we want, which, of course, leaves the entropy (and any other thermodynamic properties) unchanged. Second, we have surprisingly found the SCvH entropies are about (or ) higher than well-established experimental data at room conditions [46, 47], as mentioned in the main text. As shown in Figure S7, this discrepancy between the two sets of data is fairly uniform across a broad region of (low) temperatures and densities, including the standard condition (1 bar, 298.15K) that has been marked explicitly.
@@
To get a feeling of the impact such an entropy error can have on the interpolation, Figure S8 shows the resulting EOS using the original (but incorrect) SCvH entropy data. Note that although we have selected an DFT energy offset to connect the free energy isotherms on both sides as smoothly as possible (see panel (c)), there still remains a noticeable mismatch. On the other hand, Fig. S8(d) shows the interpolated pressure is also apparently problematic. In particular, the wild wiggling behavior at the intermediate densities can even cause violation of the basic mechanical stability constraint [21] and hence definitely unphysical.
Notice that in Fig. S8(c), the free energy isotherms of SCvH appear to be wider than the ab initio region. Recalling the basic relation , the problematic interpolation can then be remedied by globally reducing the SCvH entropy to match the correct experimental value. This leads to the final EOS shown in Fig. 3 of the main text. Note we have also added an extra entropy offset of on the DFT side to make the interpolation look slightly better. Nevertheless, this residual error is much smaller than the SCvH entropy error discussed above and does not affect our essential arguments on the accuracy and reliability of the new EOS construction approach.
@@
In principle, the remaining small error near the interpolation boundaries can be eliminated by using more accurate chemical models at low densities or electron structure methods (such as QMC) at high densities. Besides, the nuclear quantum effect may also have a slight impact on the ab initio entropy at low temperatures, which has been ignored in this work. This can be remedied either by adding some semi-empirical corrections [78, 79] or resorting to more accurate (and also more costly) path integral MD simulations [80].
@@ Appendix G More details of comparison between different Jupiter’s adiabats
Figure S9 shows some more details of Jupiter’s adiabats calculated using various hydrogen EOSs. Compared to Fig. 4 of the main text, there is additionally the adiabat derived from the CMS19 EOS [11] (yellow line), which is clearly smoother in the ab initio region than both our result (red dashed line) and MH13 (blue line). This difference originates from the significant, non-monotonic slope change of our entropy isotherms, as shown in Fig. 1(a) of the main text. In Fig. S9, we also directly illustrate the entropy of our EOS by colored regions. The change of slope is again clearly evident, indicating that our EOS better captures the pressure-induced molecular dissociation than CMS19.
By definition, our calculated adiabat (red dashed line) follows the shape of a certain “stripe” of constant entropy, as clearly shown in Fig. S9. Note that the slightly imperfect behavior of our adiabat between 1 and 10GPa corresponds to the interpolation region of our EOS; see the previous section for more discussions. However, it is worth emphasizing that this remaining error will not affect the accuracy and reliability of our adiabat at higher pressures, where the disagreements between various adiabats make the biggest difference on resulting planetary model predictions [4, 7].
In Fig. S9, our adiabat is confined to the ab initio region explored in this work, specifically up to K and . It is in principle just a matter of additional DFT-MD simulations or interpolations with other theories [21, 81, 82, 83] to extend the hydrogen EOS to higher temperatures and densities. This will enable us to accurately study Jupiter’s deeper interior beyond the red dashed line in Fig. S9, or other astrophysical objects under even more extreme conditions, such as solar-type stars.

### v2

@@ Appendix E Effect of thermodynamic inconsistency on entropy calculation
The subtleties involved in the implementation of TI arise not only from different choices of integration path, but also the reference point. In fact, the latter can be seen as an alternative manifestation of the former, which can be understood by noting that an integration path can belong to different types (i.e., isotherm or isochore first) when viewed by different reference points. As an illustration, Fig. S5 also shows the entropy reported by Miguel et al. Miguel, Y. et al. (2016) (dashed lines). According to the authors, they processed the original REOS3 data by integrating along isotherm first, hence in accordance with the red solid lines of our result. However, these two sets of entropies clearly differ by an offset in the ab initio region of the main panel. The remedy of this apparent disagreement turns out to be surprisingly easy: notice all the entropy isochores already align well, we just need to locate the reference point on another “correct” isotherm, say, 15000K. After doing this, the red solid lines then coincide perfectly with the dashed lines from Ref. Miguel, Y. et al. (2016), as shown in Figure S6.
Our reproduction of the entropy results by Miguel et al. Miguel, Y. et al. (2016) based on an ad hoc choice of the integration path and reference point, as done above, demonstrates how fragile the “global TI” approach can be for an EOS with significant thermodynamic inconsistencies. Notice the solid and dashed lines in Fig. S5 differ by roughly , which can make a big difference on planetary model predictions as shown in the main text.
@@ Appendix F More details on the thermodynamically consistent construction of our EOS
To construct an EOS that is thermodynamically consistent across several distinct theories, such as ab initio DFT and SCvH, one needs to first calculate the free energy and entropy of each region separately using the standard TI approach. We then perform a two-dimensional spline interpolation over each set of free energy data, following a similar approach to Militzer and Hubbard Militzer and Hubbard (2013). This in principle allows us to reproduce the original pressure and energy data by performing partial derivatives. The quality of such reproduction is, of course, closely related to the original level of thermodynamic consistency in each region. Once the free energy interpolation for each separate region is ready, we can then glue them together by performing another spline interpolation along their boundaries. Note one can utilize relevant partial derivative values derived from the free energy interpolations on both sides to ensure a smoother connection.
@@
In principle, the free energy interpolation would be perfectly smooth if the absolute entropy and energy (measured relative to the same baseline) were exact on both sides. However, this is not the case in practice. First, it is generally believed that only the relative energies produced by DFT (using any given exchange-correlation functional) are reliable, rather than the absolute ones. This indicates that we are allowed to globally shift the DFT energies as we want, which, of course, leaves the entropy (and any other thermodynamic properties) unchanged. Second, note that in the DFT-MD region, we have implicitly neglected the entropy contribution from the proton spin degrees of freedom, as they are irrelevant for essentially classical protons at high temperatures. However, this constant plays an important role when trying to connect with SCvH, which has taken account of accurate energy levels arising from both proton spin isomers (i.e., para- and ortho-hydrogen) when constructing the partition function of molecular hydrogen gas at low densities Saumon and Chabrier (1991). As a result, the constant must be subtracted from the original SCvH data to ensure a truly fair matching of the absolute entropies on both sides. This is precisely what we have done to produce the final EOS shown in Fig. 3 of the main text.
@@
To get a feeling on how good is the interpolation quality, it is instructive and interesting to see how bad it will become if we just mindlessly adopt the original SCvH entropy data without subtracting the constant mentioned above. The resulting EOS based on the same interpolation procedure is shown in Figure S7. Notice that although we have selected a DFT energy offset to connect the free energy isotherms on both sides as smoothly as possible (see panel (c)), there still remains a noticeable mismatch. In fact, according to the basic relation , the free energy isotherms of SCvH are apparently wider than the ab initio region due to too large entropy on the SCvH side. On the other hand, Fig. S7(d) shows the interpolated pressure is also problematic. In particular, the wild wiggling behavior at the intermediate densities can even cause violation of the basic mechanical stability constraint Saumon et al. (1995) and hence definitely unphysical.
The observations above indicate that the interpolation quality serves as a very sensitive and stringent “detector” of the accuracy of both theories. In particular, we have been forced to identify, confirm and “fix” the constant shift in original SCvH entropy data, even if we were totally unaware of this subtlety in the early stages of this work. This is a very compelling demonstration of the power of our new EOS construction protocol. In particular, such an iteratively “self-correcting” mechanism possessed by the protocol is exactly what we need to ensure steady convergence towards a conclusive hydrogen EOS for planetary modeling.
@@
Although we have achieved a fairly excellent matching between SCvH and DFT-PBE-MD, there still remains some small errors. In particular, note we actually have also added an extra (and very small) entropy offset of on the DFT side to make the interpolation shown in Fig. 3 of the main text look slightly better. In principle, such remaining errors can be eliminated by using more accurate chemical models at low densities or electron structure methods (such as QMC) at high densities. Besides, the nuclear quantum effect may also have a slight impact on the ab initio entropy at low temperatures, which has been ignored in this work. This can be remedied either by adding some semi-empirical corrections Militzer (2013); Becker et al. (2013) or resorting to more accurate (and also more costly) path integral MD simulations Morales et al. (2013).
@@ Appendix G More details of comparison between different Jupiter’s adiabats
Figure S8 shows some more details of Jupiter’s adiabats calculated using various hydrogen EOSs. Compared to Fig. 4 of the main text, there is additionally the adiabat derived from the CMS19 EOS Chabrier et al. (2019) (yellow line), which is clearly smoother in the ab initio region than both our result (red dashed line) and MH13 (blue line). This difference originates from the significant, non-monotonic slope change of our entropy isotherms, as shown in Fig. 1(a) of the main text. In Fig. S8, we also directly illustrate the entropy of our EOS by colored regions. The change of slope is again clearly evident, indicating that our EOS better captures the pressure-induced molecular dissociation than CMS19.
By definition, our calculated adiabat (red dashed line) follows the shape of a certain “stripe” of constant entropy, as clearly shown in Fig. S8. Note that the slightly imperfect behavior of our adiabat between 1 and 10GPa corresponds to the interpolation region of our EOS; see the previous section for more discussions. However, it is worth emphasizing that this remaining error will not affect the accuracy and reliability of our adiabat at higher pressures, where the disagreements between various adiabats make the biggest difference on resulting planetary model predictions Miguel, Y. et al. (2016); Howard, S. et al. (2023).
In Fig. S8, our adiabat is confined to the ab initio region explored in this work, specifically up to K and . It is in principle just a matter of additional DFT-MD simulations or interpolations with other theories Saumon et al. (1995); Chabrier and Potekhin (1998); Militzer and Ceperley (2000, 2001) to extend the hydrogen EOS to higher temperatures and densities. This will enable us to accurately study Jupiter’s deeper interior beyond the red dashed line in Fig. S8, or other astrophysical objects under even more extreme conditions, such as solar-type stars.